# A classification of Fourier summation formulas and crystalline measures

### v1

@@ 5. Almost Periodic Functions in $\mathbb{C}^{+}$
we conclude that ${\rm spec}(\Phi\circ h)\subset\{0\}\cup[b,\infty)$. If in addition ${\rm spec}(h)$ is locally finite, then since ${\rm spec}(\Phi\circ h)\cap[0,T]\subset\cup_{0\leq n\leq T/b}\big{(}\oplus^{n}{\rm spec}(h)\cap[0,T]\big{)}$, we conclude that ${\rm spec}(\Phi\circ h)$ is locally finite. Considering $\Phi(z)=e^{-iz}$ and $h(z)=g(z)-\mathbb{E}g(0)$, we deduce that $F(z)=\Phi(h(z))e^{-i\mathbb{E}g(0)-ipz}$ belongs to $\mathrm{AP}(\mathbb{C}^{+})$ and that ${\rm spec}(F)\subset\{-p/(2\pi)\}\cup[-p/(2\pi)+b,\infty)$. Moreover, ${\rm spec}(F)$ is locally finite if ${\rm spec}(f)$ also is. ∎
We also have approximation by trigonometric polynomials (see [7, p. 148])
@@ 6. Representations of Analytic Functions
Now we show $(2)\Rightarrow(3)$. Since $g$ has no zeros, Nevalinnas’s factorization for functions of bounded type [9, Theprem 9] implies that there exists a unique real-valued locally finite measure $\mu$, with $\deg(\mu)\leq 2$, and some $c,h\in\mathbb{R}$ such that
@@
belongs to AP(+)normal-$\mathrm{AP}(\mathbb{C}^{+})$ and $\mathbb{E}F(\lambda)={\boldsymbol{1}}_{\lambda\geq b}\mathbb{E}f(\lambda)$ for $\lambda\neq 0$.
@@
and, since ${\rm spec}(f)\cap(0,b)=\emptyset$ for some $b>0$, we obtain ${\rm spec}(F^{\prime})\subset[b,\infty)$. We can now apply [7, Thm. 9, p. 152] to conclude that $F\in\mathrm{AP}(\mathbb{C}^{+})$. ∎
@@
Observe that since $f$ has spectrum bounded from below, then Lemma 11(i) shows that for $m=\min\{\inf{\rm spec}(f),\inf{\rm spec}(f^{*})\}$ we have that $|f(z)e^{-2\pi imz}|+|f^{*}(z)e^{-2\pi imz}|$ is bounded for ${\rm Im}\,z>c$, for any $c>0$. This in particular shows that $f(z)e^{-2\pi|m|z}$ is bounded on the sides of any sector $S_{c}=\{z=x+iy:x>|y|/c\}$. Since $f$ has finite order, one can apply Phranmën-Lindelöf for sectors (with small $c>0$) [27, Thm. 1, p. 37] to conclude that $f(z)e^{-2\pi|m|z}$ is bounded in $S_{c}$. We then apply the same argument replacing $f(z)$ by $f^{*}(-z)$ to conclude that $f^{*}(-z)e^{-2\pi|m|z}$ is bounded in $S_{c}$. All these imply that $f$ has finite exponential type, that is,
@@
for some $C>0$. However, since $f(z+i)e^{-2\pi im(z+i)}$ is bounded on the real line, another application of Phranmën-Lindelöf [27, Thm. 3, p. 38] shows that $|f(z+i)e^{-2\pi im(z+i)}|\ll e^{C^{\prime}|y|}$ for some $C^{\prime}>0$, and so $|f(z)|\ll e^{C^{\prime\prime}|y|}$ for some $C^{\prime\prime}>0$. In particular $f$ is bounded on every horizontal strip and, by the same result, we can take $C^{\prime}=\tau$. A straightforward calculation using Cauchy’s integral formula shows that
@@
since $f$ is now bounded on horizontal strips. In particular, ${\rm spec}(f)=-{\rm spec}(f^{*})$. A well-known result of Krein [20] (see also [27, Thm. 1, p. 115]) shows that $f,f^{*}\in{\rm BT}(\mathbb{C}^{+})$ and that $\tau=\max\{\vartheta(f^{*}),\vartheta(f)\}$. Since $g(z)=f(z)e^{i\tau z}$ is bounded in $\mathbb{C}^{+}$, Lemma 11(i) shows that ${\rm spec}(f)+\tau/(2\pi)={\rm spec}(g)\subset[0,\infty)$, and so ${\rm spec}(f)\subset[-\tau/(2\pi),\infty)$. The same argument shows that ${\rm spec}(f^{*})\subset[-\tau/(2\pi),\infty)$, but since $\mathbb{E}f(\lambda)=\overline{\mathbb{E}(f^{*})(-\lambda)}$, we obtain that ${\rm spec}(f)\subset[-\tau/(2\pi),\tau/(2\pi)]$. Finally, if ${\rm spec}(f)$ is locally finite, Proposition 13 proves straightforwardly that $f$ is a trigonometric polynomial. ∎
@@ 7. Proof of Theorem 1
Let $\varphi\in C_{c}^{\infty}(\mathbb{R})$, $\int\varphi=1$, $\varphi\geq 0$, $\operatorname{\mathrm{supp}}\varphi\subset(-1,1)$ and define $\varphi_{\varepsilon}(x)=\varphi(x/\varepsilon)/\varepsilon$ for $10<\varepsilon<$ 1. Let $w,z\in R_{c}$, $T>2/c$ and
@@
is independent of $y>c_{1}$ and is nonzero only for countable many real $\lambda$’s. By condition $({\rm IV})$, we can apply Lemma 14 to deduce there is a real-valued locally finite measure $\mu$ of degree at most $2$ such that (17) is true. In particular,
@@
As in Lemma 14 one can then write $f=f_{1}-f_{2}$ such that ${\rm Re}\,f_{j}\geq 0$ in $\mathbb{C}^{+}$ and $f_{j}(i)=\int_{\mathbb{R}}\frac{\mathrm{d}\mu_{j}(t)}{2\pi(1+t^{2})}$. We can then apply Harnack’s inequality [7, p. 136], that translates to
@@
and so one can now apply the Phranmën-Lindelöf principle (as in [9, Thm. 1]) to obtain that $f(z+ic)$ is bounded in $\mathbb{C}^{+}$ for every $c>c_{1}$. Lemma 11(i) shows that $\mathbb{E}f(\lambda)=0$ for $\lambda<0$. By condition $({\rm IV})$, there is $B_{M}>0$ and $c_{2}\geq c_{1}$ (with equality only if $f(\cdot+ic_{1})\in{\rm AP}(\mathbb{R})$) such that
whenever $p(x)=\sum_{\theta\in S}p_{\theta}e^{2\pi i\theta x}$ for some finite set $S\subset[0,M]$. We conclude that
$\sum_{0\leq\lambda\leq M}|\mathbb{E}f(\lambda)|\leq B_{M}e^{2\pi c_{2}M}$
for every $M>0$, so the function $a(\lambda)$ defined in (4) is locally summable. Finally, let us prove that $(\mu,a)$ is a $\mathrm{FS}$-pair. For $z\in\mathbb{C}^{+}$ and $t\in\mathbb{R}$ let $P_{z}(t)=\frac{z}{\pi i(t^{2}-z^{2})}$. Since identity (17) holds true, a simple computation shows
@@
for all $s\in\mathbb{R}$. By Bochner’s approximation on ${\rm AP}(\mathbb{R})$, for every $h>0$, one can find a sequence of functions $d_{n}:\mathbb{R}\to[0,1)$, each with finite support and such that $\lim_{n}d_{n}(\lambda)={\boldsymbol{\delta}}_{{\rm spec}(f)}$, and that
@@
Appendix: $\mathrm{FS}$-pairs, an account of various examples.
Throughout this section we define certain families of $\mathrm{FS}$-pairs, which will be closed under scaling, translation, modulation and multiplication by scalar.
Finite Support
Let $(\mu,a)$ be a $\mathrm{FS}$-pair such that $a(\cdot)$ has finite support, this is, there is $\{p_{n}\}_{n=1}^{N}\subset\mathbb{C}$ and $\{\lambda_{n}\}_{n=1}^{N}\subset\mathbb{R}$ such that
@@
for all $\varphi\in C^{\infty}_{c}(\mathbb{R})$. It is easy to see that we must have $\mathrm{d}\mu(t)=\sum_{n=1}^{N}p_{n}e^{2\pi i\lambda_{n}t}\mathrm{d}t$, since the difference of these two measures would vanish identically over $C^{\infty}_{c}(\mathbb{R})$.
Uniformly Discrete Poisson Summation (${\rm UDPS}$)
We let the class ${\rm UDPS}$ denote the set of pairs $(\mu,a)$ such that there is $\alpha>0$, $\{\theta_{j}\}_{j=1}^{N}\subset[0,\alpha)$ and trigonometric polynomials $\{Q_{j}\}_{j=1}^{N}$ such that
@@
It is easy to see that if $(\mu,a)\in{\rm UDPS}$ then $a(\cdot)$ is also of the above form. The breakthrough of Lev & Olveskii [25] shows that $(\mu,a)\in{\rm UDPS}$ if and only if $\mu$ and $a(\cdot)$ have uniformly discrete support (assuming both are strongly tempered). We observe that we can always assume that all $Q_{j}^{\prime}s$ are equal since by classical interpolation results (for instance, Lagrange interpolation via Chebyshev polynomials) one can always find trigonometric polynomials $P_{j}$, which are $\alpha$-periodic, and such that $P_{j}(\theta_{k})={\boldsymbol{\delta}}_{j,k}$. Thus, if we let $Q=\sum_{j=1}^{N}P_{j}Q_{j}$ then $Q(\alpha n+\theta_{j})=Q_{j}(\alpha n+\theta_{j})$ for all $n\in\mathbb{Z}$, and so
@@
Real Rooted Trigonometric Polynomial ($\mathrm{RRTP}$)
@@
where $Q_{j}$ are trigonometric polynomials, $\varphi_{j}$ is the phase function associated with $E_{j}$, defined by the condition that $E_{j}(x)e^{i\varphi_{j}(x)}$ is real for all $x\in\mathbb{R}$ ($\varphi_{j}$ is uniquely defined modulo $\pi$, see (6)). We observe that if we write $E_{j}=A_{j}-iB_{j}$, were $A_{j},B_{j}$ are trigonometric polynomials which are real on the real line, then both $A,B$ have only real zeros, $A/B$ have only simple real zeros and simple poles which interlace,
$\{\gamma\in\mathbb{R}:\,\varphi_{j}(\gamma)\equiv 0\,({\rm mod}\,\pi)\}={\rm Zeros}(B_{j}/A_{j})\quad\text{and}\quad\frac{1}{\varphi_{j}^{\prime}(\gamma)}=\mathop{\mathrm{Res}}_{z=\gamma}\,(A_{j}/B_{j})>0,$
By Theorem 3 we conclude that $(\mu,\widehat{\mu})\in\rm CM$ (see below). By Remark 5, $\mathrm{RRTP}$ contains the pairs $(\mu,\widehat{\mu})$ where
@@
$Q,P$ are trigonometric polynomials, $P$ with only real roots and $m(\gamma)$ is the multiplicity of the zero $\gamma$. There measure were first introduced and studied by Kurasov and Sarnak [21]. Observe that ${\rm UDPS}\subset\mathrm{RRTP}$ since $\cup_{j}\alpha\mathbb{Z}+\theta_{j}=\text{Zeros}(\prod_{j}\sin(\pi(z-\theta_{j})/\alpha))$. Indeed $\mathrm{RRTP}$ is a much richer class since
@@
has only real roots if $U$ is a $n\times n$ unitary matrix and $l_{1},....,l_{n}\in\mathbb{R}$. The zeros of such polynomials are generically not contained in a finite union of arithmetic progressions [2]. A simple example is $P(x)=\sin(x)+\frac{1}{10}\sin(\sqrt{2}x)$. Trigonometric polynomials with only real zeros have been classified recently in [1] as restrictions of Lee-Yang polynomials. A polynomial $P:\mathbb{C}^{n}\to\mathbb{C}$ is Lee-Yang if $P(z_{1},...,z_{n})\neq 0$ whenever $\min\{\max\{|z_{1}|,...,|z_{n}|\},\max\{|z_{1}|^{-1},...,|z_{n}|^{-1}\}\}<1$. Indeed, the main result in [1] states that if $Q$ is a trigonometric polynomial with only real zeros then there exists a Lee-Yang polynomial $P(z_{1},...,z_{n})$ and positive reals $l_{1},...,l_{n}$, linear independent over $\mathbb{Q}$, such that
$Q(x)=P(e^{il_{1}x},...,e^{il_{n}x}).$
Crystalline Measures ($\rm CM$)
We let $\rm CM$ be the class of $\mathrm{FS}$-pairs $(\mu,a)$ such that both $\mu$ and $a(\cdot)$ have locally finite support. It is not hard to see that $\mathrm{RRTP}\subset\rm CM$. One main theme investigated in [21, 2, 1] is to find pairs $(\mu,a)\in\rm CM$ such that $\operatorname{\mathrm{supp}}(a)$ and/or $\operatorname{\mathrm{supp}}(\mu)$ intersects any infinite arithmetic progression at most finitely [^16] So very non-periodic. Guinand [18, p. 265], in 1959, was the first [^17] In the sense he stated in terms of a generalized Poisson summation type formula. Although his construction was selected from a book of Landau from 1909. to produce such pair. He constructs a (self-dual) $\mathrm{FS}$-pair $(\mu,\mu)$ with

### v3

@@ 5. Almost Periodic Functions in $\mathbb{C}^{+}$
we conclude that ${\rm spec}(\Phi\circ h)\subset\{0\}\cup[b,\infty)$. If in addition ${\rm spec}(h)$ is locally finite, since
${\rm spec}(\Phi\circ h)\cap[0,T]\subset\cup_{0\leq n\leq T/b}\big(\oplus^{n}{\rm spec}(h)\cap[0,T]\big),$
we conclude that ${\rm spec}(\Phi\circ h)$ is locally finite. Considering $\Phi(z)=e^{-iz}$ and $h(z)=g(z)-\mathbb{E}g(0)$, we deduce that $F(z)=\Phi(h(z))e^{-i\mathbb{E}g(0)-ipz}$ belongs to $\mathrm{AP}(\mathbb{C}^{+})$ and that ${\rm spec}(F)\subset\{-p/(2\pi)\}\cup[-p/(2\pi)+b,\infty)$. Moreover, ${\rm spec}(F)$ is locally finite if ${\rm spec}(f)$ also is. ∎
We also have approximation by trigonometric polynomials (see [10, p. 148])
@@ 6. Representations of Analytic Functions
Now we show $(2)\Rightarrow(3)$. Since $g$ has no zeros, Nevalinnas’s factorization for functions of bounded type [13, Theprem 9] implies that there exists a unique real-valued locally finite measure $\mu$, with $\deg(\mu)\leq 2$, and some $c,h\in\mathbb{R}$ such that
@@
belongs to $\mathrm{AP}(\mathbb{C}^{+})$ and $\mathbb{E}F(\lambda)={\boldsymbol{1}}_{\lambda\geq b}\mathbb{E}f(\lambda)$ for $\lambda\neq 0$.
@@
and, since ${\rm spec}(f)\cap(0,b)=\emptyset$ for some $b>0$, we obtain ${\rm spec}(F^{\prime})\subset[b,\infty)$. We can now apply [10, Thm. 9, p. 152] to conclude that $F\in\mathrm{AP}(\mathbb{C}^{+})$. ∎
@@
Observe that since $f$ has spectrum bounded from below, then Lemma 12(i) shows that for $m=\min\{\inf{\rm spec}(f),\inf{\rm spec}(f^{*})\}$ we have that $|f(z)e^{-2\pi imz}|+|f^{*}(z)e^{-2\pi imz}|$ is bounded for ${\rm Im}\,z>c$, for any $c>0$. This in particular shows that $f(z)e^{-2\pi|m|z}$ is bounded on the sides of any sector $S_{c}=\{z=x+iy:x>|y|/c\}$. Since $f$ has finite order, one can apply Phranmën-Lindelöf for sectors (with small $c>0$) [33, Thm. 1, p. 37] to conclude that $f(z)e^{-2\pi|m|z}$ is bounded in $S_{c}$. We then apply the same argument replacing $f(z)$ by $f^{*}(-z)$ to conclude that $f^{*}(-z)e^{-2\pi|m|z}$ is bounded in $S_{c}$. All these imply that $f$ has finite exponential type, that is,
@@
for some $C>0$. However, since $f(z+i)e^{-2\pi im(z+i)}$ is bounded on the real line, another application of Phranmën-Lindelöf [33, Thm. 3, p. 38] shows that $|f(z+i)e^{-2\pi im(z+i)}|\ll e^{C^{\prime}|y|}$ for some $C^{\prime}>0$, and so $|f(z)|\ll e^{C^{\prime\prime}|y|}$ for some $C^{\prime\prime}>0$. In particular $f$ is bounded on every horizontal strip and, by the same result, we can take $C^{\prime}=\tau$. A straightforward calculation using Cauchy’s integral formula shows that
@@
since $f$ is now bounded on horizontal strips. In particular, ${\rm spec}(f)=-{\rm spec}(f^{*})$. A well-known result of Krein [25] (see also [33, Thm. 1, p. 115]) shows that $f,f^{*}\in{\rm BT}(\mathbb{C}^{+})$ and that $\tau=\max\{\vartheta(f^{*}),\vartheta(f)\}$. Since $g(z)=f(z)e^{i\tau z}$ is bounded in $\mathbb{C}^{+}$, Lemma 12(i) shows that ${\rm spec}(f)+\tau/(2\pi)={\rm spec}(g)\subset[0,\infty)$, hence ${\rm spec}(f)\subset[-\tau/(2\pi),\infty)$. The same argument shows that ${\rm spec}(f^{*})\subset[-\tau/(2\pi),\infty)$, but since $\mathbb{E}f(\lambda)=\overline{\mathbb{E}(f^{*})(-\lambda)}$, we obtain that ${\rm spec}(f)\subset[-\tau/(2\pi),\tau/(2\pi)]$. Finally, if ${\rm spec}(f)$ is locally finite, Proposition 14 proves straightforwardly that $f$ is a trigonometric polynomial. ∎
@@ 7. Proof of Theorem 1
Let $\varphi\in C_{c}^{\infty}(\mathbb{R})$, $\int\varphi=1$, $\varphi\geq 0$, $\operatorname{\mathrm{supp}}\varphi\subset(-1,1)$ and define $\varphi_{\varepsilon}(x)=\varphi(x/\varepsilon)/\varepsilon$ for $0<\varepsilon<1$. Let $w,z\in R_{c}$, $T>2/c$ and
@@
is independent of $y>c_{1}$ and is nonzero only for countable many real $\lambda$’s. By condition $({\rm III}^{*})$, there is $B_{M}>0$ and $c_{2}\geq c_{1}$ (with equality only if $f(\cdot+ic_{1})\in{\rm AP}(\mathbb{R})$) such that
whenever $p(x)=\sum_{\theta\in S}p_{\theta}e^{2\pi i\theta x}$ for some finite set $S\subset[0,M]$. We conclude that ${\rm spec}(f)\subset[0,\infty)$ and
$\sum_{0\leq\lambda\leq M}|\mathbb{E}f(\lambda)|\leq B_{M}e^{2\pi c_{2}M}$
for every $M>0$, so the function $a(\lambda)$ defined in (4) is locally summable. Lemma 12(i) $f$ is bounded for ${\rm Im}\,z>2c_{1}$, in particular $\lim_{y\to\infty}{\rm Re}\,f(iy)/y=0$. By condition $({\rm IV})$, we can now apply Lemma 15 to deduce there is a real-valued locally finite measure $\mu$ of degree at most $2$ such that (17) is true. In particular,
@@
As in Lemma 15 one can then write $f=f_{1}-f_{2}$ such that ${\rm Re}\,f_{j}\geq 0$ in $\mathbb{C}^{+}$ and $f_{j}(i)=\int_{\mathbb{R}}\frac{\mathrm{d}\mu_{j}(t)}{2\pi(1+t^{2})}$. We can then apply Harnack’s inequality [10, p. 136], that translates to
@@
and so one can now apply the Phranmën-Lindelöf principle (as in [13, Thm. 1]) to obtain that $f(z+ic)$ is bounded in $\mathbb{C}^{+}$ for every $c>c_{1}$. Lemma 12(i) shows that $\mathbb{E}f(\lambda)=0$ for $\lambda<0$. Finally, let us prove that $(\mu,a)$ is a $\mathrm{FS}$-pair. For $z\in\mathbb{C}^{+}$ and $t\in\mathbb{R}$ let $P_{z}(t)=\frac{z}{\pi i(t^{2}-z^{2})}$. Since identity (17) holds true, a simple computation shows
@@
for all $s\in\mathbb{R}$. By Bochner’s approximation on ${\rm AP}(\mathbb{R})$, for every $h>0$, one can find a sequence of functions $d_{n}:\mathbb{R}\to[0,1)$, each with finite support and such that $\lim_{n}d_{n}(\lambda)={\bf 1}_{{\rm spec}(f)}(\lambda)$, and that
@@
8. $\mathrm{FS}$-pairs, an account of various examples.
Throughout this section we will define certain families of $\mathrm{FS}$-pairs for which we will try to provide as many examples as possible, old and new. We define these families so that they are closed under scaling, translation, modulation and multiplication by scalar.
8.1. Finite Support
Let $(\mu,a)$ be a $\mathrm{FS}$-pair such that $a(\cdot)$ has finite support, that is, there is $\{p_{n}\}_{n=1}^{N}\subset\mathbb{C}$ and $\{\lambda_{n}\}_{n=1}^{N}\subset\mathbb{R}$ such that
@@ 8.1. Finite Support
for all $\varphi\in C^{\infty}_{c}(\mathbb{R})$. It is easy to see that we must have $\mathrm{d}\mu(t)=\sum_{n=1}^{N}p_{n}e^{2\pi i\lambda_{n}t}\mathrm{d}t$.
8.2. Uniformly Discrete Poisson Summation ($\rm UDPS$})
We denote $\rm UDPS$} the set of $\mathrm{FS}$-pairs $(\mu,a)$ such that there is $\alpha>0$, $\{\theta_{j}\}_{j=1}^{N}\subset[0,\alpha)$ and trig-polynomials $\{Q_{j}\}_{j=1}^{N}$ where
@@ 8.2. Uniformly Discrete Poisson Summation ($\rm UDPS$})
It is easy to see that if $(\mu,a)\in{\rm UDPS}$ then $a(\cdot)$ is also of the above form. The breakthrough of Lev & Olveskii [31] shows that $(\mu,a)\in{\rm UDPS}$ if and only if $\mu$ and $a(\cdot)$ have uniformly discrete support (assuming apriori that both are strongly tempered), which justifies the name of this family. We observe that we can always assume that all $Q_{j}^{\prime}s$ are equal since by classical interpolation results (for instance, Lagrange interpolation via Chebyshev polynomials) one can always find trigonometric polynomials $P_{j}$, which are $\alpha$-periodic, and such that $P_{j}(\theta_{k})={\boldsymbol{\delta}}_{j,k}$. Thus, if we let $Q=\sum_{j=1}^{N}P_{j}Q_{j}$ then $Q(\alpha n+\theta_{j})=Q_{j}(\alpha n+\theta_{j})$ for all $n\in\mathbb{Z}$, and so
@@
8.3. Real Rooted Trigonometric Polynomial ($\mathrm{RRTP}$)
@@
where $Q_{j}$ are trigonometric polynomials, $\varphi_{j}$ is the phase function associated with $E_{j}$, defined by the condition that $E_{j}(x)e^{i\varphi_{j}(x)}$ is real for all $x\in\mathbb{R}$ ($\varphi_{j}$ is uniquely defined modulo $\pi$, see (5)). We observe that if we write $E_{j}=A_{j}-iB_{j}$, were $A_{j},B_{j}$ are trigonometric polynomials which are real on the real line, then both $A$ and $B$ have only real zeros, $A/B$ have only simple real zeros and simple poles which interlace,
[…]
@@
and $Q$ and $P$ are any given trigonometric polynomials, $P$ with only real roots and $m(\gamma)$ is the multiplicity of the zero $\gamma$. These measures were first introduced and studied by Kurasov and Sarnak [26]. Observe that ${\rm UDPS}\subset\mathrm{RRTP}$ since $\cup_{j}\alpha\mathbb{Z}+\theta_{j}=\text{Zeros}(\prod_{j}\sin(\pi(z-\theta_{j})/\alpha))$. Indeed $\mathrm{RRTP}$ is a much richer class than $\rm UDPS$} since, for instance,
@@
has only real roots if $U$ is a $n\times n$ unitary matrix and $l_{1},....,l_{n}\in\mathbb{R}$. The zeros of such polynomials are generically not contained in a finite union of arithmetic progressions [2]. Trigonometric polynomials with only real zeros have been classified recently in [1] as restrictions of Lee-Yang polynomials. A polynomial $R:\mathbb{C}^{n}\to\mathbb{C}$ is Lee-Yang if $R(z_{1},...,z_{n})\neq 0$ whenever $\min\{\max\{|z_{1}|,...,|z_{n}|\},\max\{|z_{1}|^{-1},...,|z_{n}|^{-1}\}\}<1$. Indeed, the main result in [1] states that if $P$ is a trigonometric polynomial with only real zeros then there exists a Lee-Yang polynomial $R(z_{1},...,z_{n})$ and positive reals $l_{1},...,l_{n}$, linear independent over $\mathbb{Q}$, such that
$P(x)=R(e^{il_{1}x},...,e^{il_{n}x}).$
A simple example is
$R(z_{1},z_{2})=\det\left(\text{Id}_{3}-\begin{bmatrix}0&-1/3&4/3\\ 1&0&0\\ 0&2/3&1/3\end{bmatrix}\text{diag}(z_{1},z_{1},z_{2})\right)=1+z_{1}^{2}/3-z_{2}(1/3+z_{1}^{2})$
and
$P(z)=\frac{3i}{2}R(e^{\pi i(\sqrt{2}-1)z},e^{2\pi i(1+\sqrt{2})z})=\sin(2\pi z)+3\sin(2\pi\sqrt{2}z).$
Note the characteristic polynomial of the matrix above is $(x-1)(3x^{2}+2x+3)/3$, with roots $\{1,-1/3\pm i\sqrt{8}/3\}$, and so it is an unitary matrix. Hence, the polynomial $R$ above is Lee-Yang and $P$ has only real roots.
8.4. Crystalline Measures ($\rm CM$)
We let $\rm CM$ be the class of $\mathrm{FS}$-pairs $(\mu,a)$ such that both $\mu$ and $a(\cdot)$ have locally finite support. By Theorem 3 we conclude that $\mathrm{RRTP}\subset\rm CM$. One main theme investigated in [26, 2, 1] is to find pairs $(\mu,a)\in\rm CM$ such that $\operatorname{\mathrm{supp}}(a)$ and/or $\operatorname{\mathrm{supp}}(\mu)$ intersects any infinite arithmetic progression at most finitely [^18] So very non-periodic. Guinand [22, p. 265], in 1959, was “almost the first” to produce such pair, as we shall explain below. He constructs a (self-dual) $\mathrm{FS}$-pair $(\mu,\mu)$ with